OpenStudy (anonymous):

Medal is offered!! The algebraic identity y + 15 ÷ 3 + 2 =72÷9 − 1 + y can be simplified to the following equation, which we know is true because of the Commutative Property. A. y + 8 = 8 + y B. y − 8 = 8 − y C. y + 7 = 7 + y D. y − 7 = 7 − y

OpenStudy (janu16):

Cancel y on both sides \[\frac{ 15 }{ 3 }+2=\frac{ 72 }{ 9 }-1\] Simplify \[5+2=\frac{ 72 }{ 9 }-1\] Simplify \[5+2=8−1\] \[7=8-1\] \[7=7\]

OpenStudy (janu16):

@XxmathrulezXx so what chocie?

OpenStudy (anonymous):

D?

OpenStudy (janu16):

no both are postive..

OpenStudy (anonymous):

C?

OpenStudy (janu16):

yess

OpenStudy (janu16):

your chocies are confusing cause they are all over the place lol like not in order but ya C

OpenStudy (anonymous):

Thank you for explaining!!

OpenStudy (janu16):

np

OpenStudy (janu16):

its a headche to plug everythin in the equation box and try to make it pretty and clear